OpenStudy (shicarak):

will I be distributing for this equation 3(5r-6)-36

OpenStudy (justsmile_alittle):

yes

OpenStudy (shicarak):

can someone me how one more time

KjSaif (kjsaif):

Let's simplify step-by-step.

KjSaif (kjsaif):

First you distribute.

OpenStudy (shicarak):

ok

KjSaif (kjsaif):

Distribute: =(3)(5r)+(3)(−6)+ − 36 =15r+ −18 + −36

KjSaif (kjsaif):

Then you combine like terms. :)

OpenStudy (shicarak):

will I be adding or subtracting

KjSaif (kjsaif):

adding :)

KjSaif (kjsaif):

Combine Like Terms: =15r+ −18 + −36 =(15r)+(−18+−36) =15r+ −54

KjSaif (kjsaif):

Answer: =15r − 54

OpenStudy (shicarak):

ok that's what I got but I thought I was wrong

KjSaif (kjsaif):

hope this helps! :)

OpenStudy (shicarak):

thank you

alones (alones):

@KJSaif waitt you're using https://www.mathpapa.com/algebra-calculator.html?